# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 5.3.1 Relationship between Training and Alignment
A primary line of inquiry posits that SFT and RL fundamentally rely on different learning mechanisms. pmlr-v267-chu25c provide empirical evidence that SFT tends to memorize training data, leading to poor performance on out-of-distribution (OOD) tasks. In contrast, RL demonstrates superior generalization capabilities, effectively enabling the model to adapt to unseen rules in textual and visual environments.
Deeper theoretical work seeks to explain the mechanism behind RL’s superiority. swamy2025all attribute this to the “generation-verification gap”. The authors argue that in many reasoning tasks, learning a verifier is significantly easier than learning a generator. Consequently, the value of the two-stage RL process lies in using a simpler reward model to narrow the search space, effectively guiding the policy toward a subset of optimal solutions that offline cloning cannot easily identify. This perspective is further reinforced by analyzing the scalability of these methods at inference time. setlur2025scaling prove that Verifier-Based (VB) methods, such as RL or search, possess a distinct theoretical advantage over Verifier-Free (VF) methods like behavioral cloning. The study demonstrates that as test-time compute and training data increase, the performance gap between VB and VF methods widens, with VB methods achieving superior asymptotic performance. This provides a theoretical justification for the necessity of RL in alignment, particularly for reasoning-intensive tasks where verification is feasible.
Besides, recent efforts have attempted to dissolve the strict dichotomy between SFT and RL by establishing unified theoretical paradigms. shao2024deepseekmathpushinglimitsmathematical propose a unified paradigm that encompasses SFT, Rejection Sampling Fine-Tuning (RFT), DPO [rafailov2023direct], PPO [schulman2017proximal], and thus propose GRPO. By analyzing these methods under a single lens, the authors identify the key factors that drive performance across different stages. Complementing this, ren2025learning introduce a framework to analyze the learning dynamics during both SFT and alignment phases. This framework offers explanations for counterintuitive phenomena observed during the transition between stages, such as the amplification of hallucinations, suggesting that the alignment process is governed by specific dynamic laws that persist across different algorithms.
@@ 5.3.2 The Frontier of RL
While RL has become the most effective technique for aligning LLMs, the community is currently pushing the boundaries of how RL fundamentally shapes model behavior and where it can be applied beyond standard alignment. Recent work offers a dynamic view of RL process. yao2025debaterlvrreasoningcapability propose a two-stage theory of RLVR. The authors identify an initial exploitation phase where the model reinforces high-reward tokens, leading to capability shrinkage and diversity loss, followed by an exploration phase where latent, optimal low-probability tokens are boosted, expanding the capability boundary. This mechanism highlights the critical need for training strategies that can navigate the trade-off between exploitation and exploration.
@@
Complementing this, researchers are re-evaluating the specific signals used for optimization. zhu2025surprisingeffectivenessnegativereinforcement decompose learning signals into positive and negative reinforcement. The study reveals that while positive reinforcement improves greedy decoding (Pass@1), it often causes distribution collapse. However, when focusing on suppressing incorrect paths, it is surprisingly effective at maintaining diversity and improving performance across the entire Pass@k spectrum. Furthermore, the stability of RL training is being examined at the token level. yang2025letlowprobabilitytokensoverdominate identify a gradient anomaly where low-probability tokens generate disproportionately large gradient magnitudes, suppressing the learning of high-probability tokens. By proposing methods like Advantage Reweighting, this work demonstrates that balancing token-level contributions is essential for stable optimization in complex reasoning tasks.
@@
As models move toward generating longer chain-of-thought (LongCoT), the quadratic computational cost of attention becomes a bottleneck for RL training. The frontier of RL is thus exploring architecture-agnostic scaling methods. aghajohari2025markovianthinkerarchitectureagnosticlinear introduce a “Markovian Thinking” paradigm. By segmenting the reasoning process into chunks with limited state carryover, this approach achieves linear scaling with reasoning length. This allows RL to be applied to extremely long reasoning trajectories with significantly reduced computational overhead, matching or exceeding the performance of traditional full-context RL.
@@
Finally, RL is expanding from static reasoning tasks to dynamic, long-horizon agentic environments. A major challenge here is the “cold start” problem in sparse-reward settings. zhang2025agentlearningearlyexperience propose an “Early Experience” paradigm that bridges imitation learning and RL. By utilizing a model’s own exploration of future states as a self-supervised signal, agents can bootstrap learning without immediate external rewards. Simultaneously, managing context in multi-turn agent interactions is critical. tang2025turnlimitstrainingdeep address the limitation of fixed context windows in long-cycle agent tasks. The proposed “DeepMiner” framework and dynamic sliding window strategy enable agents to maintain coherent reasoning over hundreds of turns, demonstrating that by constructing complex, verifiable tasks, RL can drive agents to develop deep reasoning capabilities that transcend simple instruction following.
@@ 6.1 Fundamental Problems
These two questions, concerning the elicitation of latent behaviors through context and the dynamics of computational scaling, form the theoretical bedrock of the Inference Stage. While the former investigates how the model interprets its input to narrow down the task space, the latter focuses on how it allocates internal and external resources to navigate complex problem-solving. The landscape of current theoretical consideration in this stage is summarized in fig. 6. In what follows, we begin by reviewing the core theories and methods developed to address these challenges.
@@ 6.2.1 Prompt Engineering
Prompt engineering steers an LLM at inference time by modifying the input sequence, without updating model parameters [lester2021power, sahoo2024systematic, schulhoff2024prompt]. It serves as the main interface that translates a user intent into a form the model can follow, and it often determines whether the model uses prior knowledge, follows constraints, or produces structured outputs [brown2020language, schick2021exploiting, chen2023unleashing, sahoo2024systematic]. Beyond being a practical technique, prompt engineering also provides a window into the model’s internal behavior, since small changes in the prompt can lead to large changes in the generated distribution. Many works have begun to study prompting beyond heuristics, aiming to understand why prompt choices can reliably reshape the model’s behavior and internal computation. Based on current research, we categorize these investigations into the following four core dimensions.
@@
Prompt Design and Structured Prompting. This direction focuses on designing the prompt structure so that the model’s next-token prediction aligns with the intended task and output format. Early prompt-based formulations such as PET convert classification into cloze-style templates with label words, demonstrating that prompt design can be viewed as a task re-parameterization [schick2021exploiting]. However, prompt form can introduce large variance, zhao2021calibrate show that few-shot accuracy is highly sensitive to demonstration order and formatting, and proposes contextual calibration to correct systematic biases induced by the prompt. From a mechanism perspective, webson2022prompt find that models can succeed even when prompt semantics is weak or misleading, suggesting that surface cues and distributional patterns often dominate literal instruction understanding. Similarly, min2022rethinking report that the correctness of labels in demonstrations can be less important than specifying the input space, label space, and input-output format, which further highlights the role of prompt structure.
@@
Automated and Learnable Prompts. This direction replaces manual prompt crafting with optimization or learning procedures, aiming to systematically search for effective prompts and reduce prompt sensitivity. Early studies show that even short discrete triggers can reliably elicit target behaviors, and such prompts are often hard to interpret [shin2020autoprompt, deng2022rlprompt]. With the rise of LLMs, a recent trend treats the model itself as a prompt optimizer and improves instructions through iterative refinement based on feedback. OPRO formulates prompt search as a black-box optimization loop, where the meta-prompt records candidate instructions and their scores so that the LLM proposes better instructions in subsequent iterations [yang2023large]. Evolutionary-style methods provide another effective route, where prompts are iteratively mutated and selected to improve task performance [guo2023connecting, fernando2023promptbreeder]. To make refinement more controlled and reusable, tang2025unleashing analyze prompt updates through an analogy to gradient-based optimization and designs update rules that retrieve strong candidates while limiting the edit magnitude, while yang2024ampo organize optimization into a multi-branched prompt structure that is updated using failure cases as feedback. Rather than optimizing a single string, khattab2024dspy introduce a compiler-style framework that optimizes instructions and demonstrations for multi-stage LM pipelines under an end-task metric. Recent work also begins to systematize both practice and theory: trivedi2025align study prompt optimization for alignment-style objectives and provides principled guarantees, agrawal2025gepa propose reflective evolution that summarizes failures into reusable natural-language rules, and murthy2025promptomatix present an end-to-end framework that composes modular optimizers for automatic prompt search. For continuous prompting, hu2024fundamental characterize universality, capacity, and efficiency limits of prompt tuning in simplified Transformer settings. A systematic survey further consolidates automatic prompt optimization methods and clarifies emerging evaluation protocols and open challenges [ramnath2025systematic].
@@
Mechanisms and Diagnostic Tools for Prompting. This dimension connects prompt choices to the model’s internal computation and develops diagnostic tools that localize where and how a prompt steers generation. Mechanistic analyses identify concrete routing and copying circuits that are activated by structured context. Induction-style mechanisms provide a canonical example, where repeated patterns in a prompt trigger copying and support generalization [olsson2022context]. Subsequent studies characterize when such circuits emerge and what subcomponents are required, providing causal evidence that prompt structure can selectively activate specialized components [reddy2023mechanistic, singh2024needs, d2025selective]. Prompt influence can also be localized at the token level. feng2024unveiling propose Token Distribution Dynamics, which attributes generation to specific prompt tokens by tracking distribution dynamics over the vocabulary and enables targeted prompt edits for controlled generation. At the circuit level, recent work reduces the reliance on handcrafted analyses by introducing automated discovery and richer causal traces. conmy2023towards propose ACDC to automate circuit discovery via activation patching, while ameisen2025circuit construct attribution graphs that map information flow on individual prompts using interpretable replacement models. Complementary diagnostics infer and manipulate the prompt–computation interface at a higher abstraction. elhelo2025inferring infer attention-head functionality directly from model parameters, and representation-based steering methods extract interpretable directions from prompt contrasts to control generation [zou2023representation, turner2023steering]. At a more theoretical level, kim2025theoretical formalize prompting as varying an external program under a fixed Transformer executor, define the prompt-induced hypothesis class, and give a constructive decomposition that separates routing via attention, local arithmetic via feed-forward layers, and depth-wise composition. This formulation clarifies expressivity and makes explicit the limits imposed by prompt length and precision.
@@
Reliability, Generalization, and Security. The flexibility of the prompting interface creates significant challenges for system robustness, as adversarial inputs can systematically exploit the model’s instruction-following mechanisms to bypass safety alignment [nan2024jailbreak, li2025revisiting]. wallace2024instruction define an explicit instruction hierarchy and show that training models to follow prioritized rules improves robustness to prompt injection, including attack types not seen during training. Similarly, zhang2024defending argue that jailbreaking succeeds when helpfulness and safety goals conflict, and they reduce attack success by enforcing goal prioritization during inference and training. Mechanistic evidence further connects injection to internal routing. hung2025attention identify a distraction effect where specific attention heads shift focus from the original instruction to injected text, enabling training-free detection by monitoring these attention patterns. jiang2024robustkv show that many jailbreak prompts work by reallocating attention such that harmful tokens remain in cache, and they mitigate attacks by evicting low-importance key–value entries to suppress the concealed query signal. At the representation level, ball2024understanding extract transferable jailbreak vectors from activations and provide evidence that successful jailbreaks suppress internal “harmfulness” features, while kirch2025features reveal that jailbreak success is supported by heterogeneous and often non-linear prompt features and validate them via probe-guided latent interventions. These developments suggest that prompt engineering in deployed systems requires not only performance tuning but also principled auditing of how instruction conflicts are represented and resolved inside the forward pass [rossi2024early, yao2024survey].
@@ 6.2.2 In-Context Learning
Transformer-based Large Language Models (LLMs) [vaswani2017attention] has shown amazing in-context learning (ICL) capabilities [brown2020language, wei2022emergent, dong2022survey, liu2023pre]. ICL can be viewed as a form of few-shot learning, where the model is provided with a small number of input-label pairs as examples. Without the need for parameter updates, the model can recognize the task at hand and provide the desired answer for a given query. This fantastic capability enables pre-trained LLMs such as GPT models to be generalized in wide downstream tasks conveniently. Despite the good performance of the ICL capabilities, the mechanism of ICL still remains an open question. Many works have begun to analyze the source of ICL capabilities from different perspectives.
@@
Algorithmic Camp. This camp believes that ICL can learn the ability to execute algorithms during the pre-training phase, and then executes algorithms for different tasks during ICL inference [icl_tf_algorithms, icl_tf_linear, icl_tf_algorithm_selection]. Therefore, the algorithmic camp primarily explores the ICL mechanism through methods such as Transformer’s ability to learn certain function family. To understand how LLMs perform ICL inference without parameter updates, an intuitive idea is that there may be some implicitly updating in the model’s architecture. Following this motivation, icl_dual point out that Transformer implicitly fine-tunes during ICL inference, building upon the dual form of the attention mechanism proposed by dual_1964, dual_attention. There are also some works involves the use of the specific construction of weights, that is, assuming the parameters of the Transformer (e.g., ${\bm{W}}_{Q}$, ${\bm{W}}_{K}$, ${\bm{W}}_{V}$) have specific forms, thereby enabling the model’s forward computation to execute a certain algorithm that is easy to interpret. icl_gd_akyurek reveal that under certain constructions, Transformer can implement simple basic operations (mov, mul, div and aff), which can be combined to further perform gradient descent. icl_gd_oswald_2023 provide a more concise and appealing construction for solving least squares solutions in the linear attention setting, which is further adopted and followed by more works [icl_tf_CasualLMnotgood, icl_oswald_mesa_layer]. icl_gd_matengyu theoretically prove that when the covariates are sampled from a Gaussian distribution, the pretraining loss with a single-layer linear attention will be achieved at optimal minimization through a one-step gradient descent. Based on the construction, icl_tf_CasualLMnotgood analyzes such algorithm executed by Transformers under the casual mask setting, indicating that such construction will lead to an online gradient descent algorithm with non-decaying step size, which can not guarantee convergence to the optimal solution. Unlike CasualLM, it has been proved that PrefixLM PrefixLM can achieve theoretically optimal solutions. Similarly, icl_oswald_mesa_layer propose a new constructive approach under the auto-regressive setting and reach similar conclusions related to online gradient descent. Furthermore, it introduce mesa-layer through reverse engineering: by solving an optimization problem similar to ridge regression to output the next layer’s token representations. Furthermore, icl_tf_unstructured_data explore the ICL ability on linear regression tasks from the perspective of unstructured data, where positional encoding and multi-head attention can bring better predictive performance to ICL. The work above mostly considers the setup of linear attention. More considerations from nonlinear settings are also proposed, where the abilities of Transformer to learn a wider range of nonlinear functions are further explored [icl_gd_kernel_mit, icl_tf_softmax].
@@
Representation Camp. This camp posits that LLMs store memories about various topics during the pretraining process, and in-context learning retrieves contextually relevant topics during inference based on demonstrations. icl_bayesinference demonstrate that even in the case of a distribution mismatch, the asymptotic prediction error for in-context learning achieves optimality when the signal pertaining to the latent concept in each prompt example surpasses the error arising from the distribution mismatch. Further, they create a new small-scale synthetic dataset called the Generative IN-Context learning dataset (GINC) to study the mechanism of ICL. It is discovered that both Transformers and LSTMs have the ability to learn in-context, and this capability improves with the length and quantity of demonstrations. Similarly, icl_latent establish a general data generation process on a causal graph composed of three variables and demonstrated that the predictor can reach optimality when using latent variables to select a finite number of examples. Building upon this, they propose an efficient example selection algorithm capable of choosing examples on a smaller LLM and directly generalizing to other LLMs. icl_label_not_important conduct experiments across 12 models, including GPT-3, and find that replacing labels in the input-label pairs with random ones during ICL inference results in only marginal decreases in performance, which contrasts somewhat with the findings of icl_bayesinference. Furthermore, they identify other aspects that have a greater impact on performance, revealing that the accuracy of ICL depends on the independent specification of the input and label spaces, the distribution of the input text, and even the format of the input-output pairs. They argue that LLMs do not learn new tasks during ICL but rather use demonstrations information to locate tasks or topics, and the ability to perform tasks is learned during pretraining. icl_data_generation systematically understand existing efforts from the perspective of data generation. They categorize existing research efforts into these two learning frameworks and establish transferability between them.

### v2

@@ 5.3.1 Relationship between Training and Alignment
A primary line of inquiry posits that SFT and RL fundamentally rely on different learning mechanisms. Chu et al. (2025) provide empirical evidence that SFT tends to memorize training data, leading to poor performance on out-of-distribution (OOD) tasks. In contrast, RL demonstrates superior generalization capabilities, effectively enabling the model to adapt to unseen rules in textual and visual environments.
Deeper theoretical work seeks to explain the mechanism behind RL’s superiority. Swamy et al. (2025) attribute this to the “generation-verification gap”. The authors argue that in many reasoning tasks, learning a verifier is significantly easier than learning a generator. Consequently, the value of the two-stage RL process lies in using a simpler reward model to narrow the search space, effectively guiding the policy toward a subset of optimal solutions that offline cloning cannot easily identify. This perspective is further reinforced by analyzing the scalability of these methods at inference time. Setlur et al. (2025) prove that Verifier-Based (VB) methods, such as RL or search, possess a distinct theoretical advantage over Verifier-Free (VF) methods like behavioral cloning. The study demonstrates that as test-time compute and training data increase, the performance gap between VB and VF methods widens, with VB methods achieving superior asymptotic performance. This provides a theoretical justification for the necessity of RL in alignment, particularly for reasoning-intensive tasks where verification is feasible.
Besides, recent efforts have attempted to dissolve the strict dichotomy between SFT and RL by establishing unified theoretical paradigms. Shao et al. (2024) propose a unified paradigm that encompasses SFT, Rejection Sampling Fine-Tuning (RFT), DPO (Rafailov et al., 2023), PPO (Schulman et al., 2017), and thus propose GRPO. By analyzing these methods under a single lens, the authors identify the key factors that drive performance across different stages. Complementing this, Ren and Sutherland (2025) introduce a framework to analyze the learning dynamics during both SFT and alignment phases. This framework offers explanations for counterintuitive phenomena observed during the transition between stages, such as the amplification of hallucinations, suggesting that the alignment process is governed by specific dynamic laws that persist across different algorithms.
@@ 5.3.2 The Frontier of RL
While RL has become the most effective technique for aligning LLMs, the community is currently pushing the boundaries of how RL fundamentally shapes model behavior and where it can be applied beyond standard alignment. Recent work offers a dynamic view of RL process. Yao et al. (2025d) propose a two-stage theory of RLVR. The authors identify an initial exploitation phase where the model reinforces high-reward tokens, leading to capability shrinkage and diversity loss, followed by an exploration phase where latent, optimal low-probability tokens are boosted, expanding the capability boundary. This mechanism highlights the critical need for training strategies that can navigate the trade-off between exploitation and exploration.
@@
Complementing this, researchers are re-evaluating the specific signals used for optimization. Zhu et al. (2025f) decompose learning signals into positive and negative reinforcement. The study reveals that while positive reinforcement improves greedy decoding (Pass@1), it often causes distribution collapse. However, when focusing on suppressing incorrect paths, it is surprisingly effective at maintaining diversity and improving performance across the entire Pass@k spectrum. Furthermore, the stability of RL training is being examined at the token level. Yang et al. (2025b) identify a gradient anomaly where low-probability tokens generate disproportionately large gradient magnitudes, suppressing the learning of high-probability tokens. By proposing methods like Advantage Reweighting, this work demonstrates that balancing token-level contributions is essential for stable optimization in complex reasoning tasks.
@@
As models move toward generating longer chain-of-thought (LongCoT), the quadratic computational cost of attention becomes a bottleneck for RL training. The frontier of RL is thus exploring architecture-agnostic scaling methods. Aghajohari et al. (2025) introduce a “Markovian Thinking” paradigm. By segmenting the reasoning process into chunks with limited state carryover, this approach achieves linear scaling with reasoning length. This allows RL to be applied to extremely long reasoning trajectories with significantly reduced computational overhead, matching or exceeding the performance of traditional full-context RL.
@@
Finally, RL is expanding from static reasoning tasks to dynamic, long-horizon agentic environments. A major challenge here is the “cold start” problem in sparse-reward settings. Zhang et al. (2025b) propose an “Early Experience” paradigm that bridges imitation learning and RL. By utilizing a model’s own exploration of future states as a self-supervised signal, agents can bootstrap learning without immediate external rewards. Simultaneously, managing context in multi-turn agent interactions is critical. Tang et al. (2025b) address the limitation of fixed context windows in long-cycle agent tasks. The proposed “DeepMiner” framework and dynamic sliding window strategy enable agents to maintain coherent reasoning over hundreds of turns, demonstrating that by constructing complex, verifiable tasks, RL can drive agents to develop deep reasoning capabilities that transcend simple instruction following. More recently, Qian et al. (2026) identify the internal factors driving agent actions regardless of the task outcome and operates hierarchically to manage the complexity of agent interactions, further shifting the research focus to the agent era.
@@ 6.1 Fundamental Problems
These two questions, concerning the elicitation of latent behaviors through context and the dynamics of computational scaling, form the theoretical bedrock of the Inference Stage. While the former investigates how the model interprets its input to narrow down the task space, the latter focuses on how it allocates internal and external resources to navigate complex problem-solving. The landscape of current theoretical consideration in this stage is summarized in fig.˜6. In what follows, we begin by reviewing the core theories and methods developed to address these challenges.
@@ 6.2.1 Prompt Engineering
Prompt engineering steers an LLM at inference time by modifying the input sequence, without updating model parameters (Lester et al., 2021; Sahoo et al., 2024; Schulhoff et al., 2024). It serves as the main interface that translates a user intent into a form the model can follow, and it often determines whether the model uses prior knowledge, follows constraints, or produces structured outputs (Brown et al., 2020; Schick and Schütze, 2021; Chen et al., 2023; Sahoo et al., 2024). Beyond being a practical technique, prompt engineering also provides a window into the model’s internal behavior, since small changes in the prompt can lead to large changes in the generated distribution. Many works have begun to study prompting beyond heuristics, aiming to understand why prompt choices can reliably reshape the model’s behavior and internal computation. Based on current research, we categorize these investigations into the following four core dimensions.
@@
Prompt Design and Structured Prompting. This direction focuses on designing the prompt structure so that the model’s next-token prediction aligns with the intended task and output format. Early prompt-based formulations such as PET convert classification into cloze-style templates with label words, demonstrating that prompt design can be viewed as a task re-parameterization (Schick and Schütze, 2021). However, prompt form can introduce large variance, Zhao et al. (2021) show that few-shot accuracy is highly sensitive to demonstration order and formatting, and proposes contextual calibration to correct systematic biases induced by the prompt. From a mechanism perspective, Webson and Pavlick (2022) find that models can succeed even when prompt semantics is weak or misleading, suggesting that surface cues and distributional patterns often dominate literal instruction understanding. Similarly, Min et al. (2022b) report that the correctness of labels in demonstrations can be less important than specifying the input space, label space, and input-output format, which further highlights the role of prompt structure.
@@
Automated and Learnable Prompts. This direction replaces manual prompt crafting with optimization or learning procedures, aiming to systematically search for effective prompts and reduce prompt sensitivity. Early studies show that even short discrete triggers can reliably elicit target behaviors, and such prompts are often hard to interpret (Shin et al., 2020; Deng et al., 2022). With the rise of LLMs, a recent trend treats the model itself as a prompt optimizer and improves instructions through iterative refinement based on feedback. OPRO formulates prompt search as a black-box optimization loop, where the meta-prompt records candidate instructions and their scores so that the LLM proposes better instructions in subsequent iterations (Yang et al., 2023a). Evolutionary-style methods provide another effective route, where prompts are iteratively mutated and selected to improve task performance (Guo et al., 2023b; Fernando et al., 2023). To make refinement more controlled and reusable, Tang et al. (2025c) analyze prompt updates through an analogy to gradient-based optimization and designs update rules that retrieve strong candidates while limiting the edit magnitude, while Yang et al. (2024a) organize optimization into a multi-branched prompt structure that is updated using failure cases as feedback. Rather than optimizing a single string, Khattab et al. (2024) introduce a compiler-style framework that optimizes instructions and demonstrations for multi-stage LM pipelines under an end-task metric. Recent work also begins to systematize both practice and theory: Trivedi et al. (2025) study prompt optimization for alignment-style objectives and provides principled guarantees, Agrawal et al. (2025) propose reflective evolution that summarizes failures into reusable natural-language rules, and Murthy et al. (2025) present an end-to-end framework that composes modular optimizers for automatic prompt search. For continuous prompting, Hu et al. (2024) characterize universality, capacity, and efficiency limits of prompt tuning in simplified Transformer settings. A systematic survey further consolidates automatic prompt optimization methods and clarifies emerging evaluation protocols and open challenges (Ramnath et al., 2025).
@@
Mechanisms and Diagnostic Tools for Prompting. This dimension connects prompt choices to the model’s internal computation and develops diagnostic tools that localize where and how a prompt steers generation. Mechanistic analyses identify concrete routing and copying circuits that are activated by structured context. Induction-style mechanisms provide a canonical example, where repeated patterns in a prompt trigger copying and support generalization (Olsson et al., 2022). Subsequent studies characterize when such circuits emerge and what subcomponents are required, providing causal evidence that prompt structure can selectively activate specialized components (Reddy, 2023; Singh et al., 2024; D’Angelo et al., 2025). Prompt influence can also be localized at the token level. Feng et al. (2024b) propose Token Distribution Dynamics, which attributes generation to specific prompt tokens by tracking distribution dynamics over the vocabulary and enables targeted prompt edits for controlled generation. At the circuit level, recent work reduces the reliance on handcrafted analyses by introducing automated discovery and richer causal traces. Conmy et al. (2023) propose ACDC to automate circuit discovery via activation patching, while Ameisen et al. (2025) construct attribution graphs that map information flow on individual prompts using interpretable replacement models. Complementary diagnostics infer and manipulate the prompt–computation interface at a higher abstraction. Elhelo and Geva (2025) infer attention-head functionality directly from model parameters, and representation-based steering methods extract interpretable directions from prompt contrasts to control generation (Zou et al., 2023; Turner et al., 2023). At a more theoretical level, Kim et al. (2025b) formalize prompting as varying an external program under a fixed Transformer executor, define the prompt-induced hypothesis class, and give a constructive decomposition that separates routing via attention, local arithmetic via feed-forward layers, and depth-wise composition. This formulation clarifies expressivity and makes explicit the limits imposed by prompt length and precision.
@@
Reliability, Generalization, and Security. The flexibility of the prompting interface creates significant challenges for system robustness, as adversarial inputs can systematically exploit the model’s instruction-following mechanisms to bypass safety alignment (Nan et al., 2024; Li et al., 2025e). Wallace et al. (2024) define an explicit instruction hierarchy and show that training models to follow prioritized rules improves robustness to prompt injection, including attack types not seen during training. Similarly, Zhang et al. (2024d) argue that jailbreaking succeeds when helpfulness and safety goals conflict, and they reduce attack success by enforcing goal prioritization during inference and training. Mechanistic evidence further connects injection to internal routing. Hung et al. (2025) identify a distraction effect where specific attention heads shift focus from the original instruction to injected text, enabling training-free detection by monitoring these attention patterns. Jiang et al. (2024a) show that many jailbreak prompts work by reallocating attention such that harmful tokens remain in cache, and they mitigate attacks by evicting low-importance key–value entries to suppress the concealed query signal. At the representation level, Ball et al. (2024) extract transferable jailbreak vectors from activations and provide evidence that successful jailbreaks suppress internal “harmfulness” features, while Kirch et al. (2025) reveal that jailbreak success is supported by heterogeneous and often non-linear prompt features and validate them via probe-guided latent interventions. These developments suggest that prompt engineering in deployed systems requires not only performance tuning but also principled auditing of how instruction conflicts are represented and resolved inside the forward pass (Rossi et al., 2024; Yao et al., 2024b).
@@ 6.2.2 In-Context Learning
Transformer-based Large Language Models (LLMs) (Vaswani et al., 2017b) has shown amazing in-context learning (ICL) capabilities (Brown et al., 2020; Wei et al., 2022b; Dong et al., 2022; Liu et al., 2023b). ICL can be viewed as a form of few-shot learning, where the model is provided with a small number of input-label pairs as examples. Without the need for parameter updates, the model can recognize the task at hand and provide the desired answer for a given query. This fantastic capability enables pre-trained LLMs such as GPT models to be generalized in wide downstream tasks conveniently. Despite the good performance of the ICL capabilities, the mechanism of ICL still remains an open question. Many works have begun to analyze the source of ICL capabilities from different perspectives.
@@
Algorithmic Camp. This camp believes that ICL can learn the ability to execute algorithms during the pre-training phase, and then executes algorithms for different tasks during ICL inference (Li et al., 2023a; Zhang et al., 2023; Bai et al., 2023b). Therefore, the algorithmic camp primarily explores the ICL mechanism through methods such as Transformer’s ability to learn certain function family. To understand how LLMs perform ICL inference without parameter updates, an intuitive idea is that there may be some implicitly updating in the model’s architecture. Following this motivation, Dai et al. (2022) point out that Transformer implicitly fine-tunes during ICL inference, building upon the dual form of the attention mechanism proposed by Aiserman et al. (1964); Irie et al. (2022). There are also some works involves the use of the specific construction of weights, that is, assuming the parameters of the Transformer (e.g., ${\bm{W}}_{Q}$, ${\bm{W}}_{K}$, ${\bm{W}}_{V}$) have specific forms, thereby enabling the model’s forward computation to execute a certain algorithm that is easy to interpret. Akyürek et al. (2022) reveal that under certain constructions, Transformer can implement simple basic operations (mov, mul, div and aff), which can be combined to further perform gradient descent. Von Oswald et al. (2023a) provide a more concise and appealing construction for solving least squares solutions in the linear attention setting, which is further adopted and followed by more works (Ding et al., 2023; Von Oswald et al., 2023b). Mahankali et al. (2023) theoretically prove that when the covariates are sampled from a Gaussian distribution, the pretraining loss with a single-layer linear attention will be achieved at optimal minimization through a one-step gradient descent. Based on the construction, Ding et al. (2023) analyze such algorithm executed by Transformers under the casual mask setting, indicating that such construction will lead to an online gradient descent algorithm with non-decaying step size, which can not guarantee convergence to the optimal solution. Unlike CasualLM, it has been proved that PrefixLM (Roberts et al., 2019) can achieve theoretically optimal solutions. Similarly, Von Oswald et al. (2023b) propose a new constructive approach under the auto-regressive setting and reach similar conclusions related to online gradient descent. Furthermore, it introduce mesa-layer through reverse engineering: by solving an optimization problem similar to ridge regression to output the next layer’s token representations. Furthermore, Xing et al. (2024) explore the ICL ability on linear regression tasks from the perspective of unstructured data, where positional encoding and multi-head attention can bring better predictive performance to ICL. The work above mostly considers the setup of linear attention. More considerations from nonlinear settings are also proposed, where the abilities of Transformer to learn a wider range of nonlinear functions are further explored (Cheng et al., 2023; Collins et al., 2024).
@@
Representation Camp. This camp posits that LLMs store memories about various topics during the pretraining process, and in-context learning retrieves contextually relevant topics during inference based on demonstrations. Xie et al. (2021) demonstrate that even in the case of a distribution mismatch, the asymptotic prediction error for in-context learning achieves optimality when the signal pertaining to the latent concept in each prompt example surpasses the error arising from the distribution mismatch. Further, they create a new small-scale synthetic dataset called the Generative IN-Context learning dataset (GINC) to study the mechanism of ICL. It is discovered that both Transformers and LSTMs have the ability to learn in-context, and this capability improves with the length and quantity of demonstrations. Similarly, Wang et al. (2024c) establish a general data generation process on a causal graph composed of three variables and demonstrated that the predictor can reach optimality when using latent variables to select a finite number of examples. Building upon this, they propose an efficient example selection algorithm capable of choosing examples on a smaller LLM and directly generalizing to other LLMs. Min et al. (2022a) conduct experiments across 12 models, including GPT-3, and find that replacing labels in the input-label pairs with random ones during ICL inference results in only marginal decreases in performance, which contrasts somewhat with the findings of Xie et al. (2021). Furthermore, they identify other aspects that have a greater impact on performance, revealing that the accuracy of ICL depends on the independent specification of the input and label spaces, the distribution of the input text, and even the format of the input-output pairs. They argue that LLMs do not learn new tasks during ICL but rather use demonstrations information to locate tasks or topics, and the ability to perform tasks is learned during pretraining. Mao et al. (2024) systematically understand existing efforts from the perspective of data generation. They categorize existing research efforts into these two learning frameworks and establish transferability between them.